OpenStudy (anonymous):

find the general term \[a _{n}\] for the geometric sequence \[a_{1}=-2\] and \[a_{2}=6\]

OpenStudy (anonymous):

find the general term \[a _{n}\] for the geometric sequence \[a_{1}=-2\] and \[a_{2}=6\]

OpenStudy (zarkon):

what is 6/(-2)?

OpenStudy (anonymous):

-3 @Zarkon

OpenStudy (zarkon):

\[-2=-2\cdot (-3)^0\] \[6=-2\cdot(-3)^1\] \[\cdots\]

OpenStudy (anonymous):

I'm confused... I see how you get the 6 but what are the steps to finding the answer i need to be correct

OpenStudy (anonymous):

so the general term would be -3???

OpenStudy (zarkon):

\[a_1=-2(-3)^0=-2(-3)^{1-1}\] \[a_2=-2(-3)^1=-2(-3)^{2-1}\] \[a_3=-2(-3)^2=-2(-3)^{3-1}\] \[\vdots\] \[a_n=\cdots\]

OpenStudy (zarkon):

the common ratio is -3, but that is not the general term

OpenStudy (anonymous):

Oh I see

OpenStudy (anonymous):

well I have to go over the study plan because I have no clue what Im supposed to be doing here